OpenStudy (babyslapmafro):

Please help me find the minimum speed of the particle and its location when it has this speed.

OpenStudy (babyslapmafro):

\[r(t)=12\sqrt{t} i+t ^{\frac{ 3 }{ 2 }}j\]

OpenStudy (babyslapmafro):

t>0

OpenStudy (babyslapmafro):

i found the velocity vector and the equation for speed @ t

OpenStudy (babyslapmafro):

\[v(t)=\frac{ 6 }{ \sqrt{t} }i+\frac{ 3 }{ 2 }\sqrt{t}j\]

OpenStudy (babyslapmafro):

speed @ t: \[\sqrt{\frac{ 36 }{ t }+\frac{ 9 }{ 4 }t}\]

OpenStudy (babyslapmafro):

nvm i got it